# Isometric embeddings of surfaces for scl

### v2

@@ Acknowledgements
First and foremost, I would like to thank my supervisor Henry Wilton for our weekly discussions and his constant support. Thanks are also due to Francesco Fournier-Facio and Kevin Li for helpful conversations, as well as to the anonymous referee for careful reading and detailed comments. Financial support from an EPSRC PhD studentship is gratefully acknowledged.
@@ 2. Scl and the relative Gromov seminorm
In Definition 2.3, replacing $\mathbb{R}$ by $\mathbb{Q}$ or $\mathbb{Z}$ does not change what it means for a group homomorphism to be $\operatorname{scl}$-preserving or a (strong) isometric embedding for $\operatorname{scl}$.
@@
    with $k_{i}\in\mathbb{Z}\smallsetminus 0$. Then $\varphi$ is a strong isometric embedding for $\operatorname{scl}$.
@@ 4. Standard form for admissible surfaces
Observe first that we can harmlessly remove any disc component of $\Sigma$, since this does not change $\chi^{-}(\Sigma)$. We then say that $\Sigma$ is disc-free.
@@
In the context of $\operatorname{scl}$, it is a classical fact [cal-scl]*Proposition 2.13 that one can work with monotone admissible surfaces only. Our definition of monotonicity allows us to adapt this to the relative Gromov seminorm [^4] This would be false with the usual definition — see Footnote 3.. Calegari’s proof [cal-scl]*Proposition 2.13 works in our context, where one should deal with boundary components of $\Sigma$ mapping to distinct components of $\coprod_{i}S^{1}$ separately.
@@
We have an admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$ in $H_{2}(S,c)$, and we may assume that $f$ is monotone, transverse, and incompressible as explained in §4.a and §4.b. Since both $S$ and $\Sigma$ are oriented, and since cellular discs map homeomorphically into $S$, they can be of two types: either they preserve the orientation or they reverse it. Having cellular discs of opposite orientations is undesirable, and we are now going to modify $\Sigma$ to avoid this situation as much as possible.
@@
Assume first that there is a connected component of $\Sigma$ containing cellular discs of two different types. In particular, $\Sigma$ must contain two cellular discs of opposite types that are adjacent via a $1$-handle. Since $S$ is a surface, those two cellular discs must map to the same $2$-cell of $S$, and we are in the situation of Figure 3 (pictures are given for the case of a $2$-cell of degree $3$, but the general case is similar).
@@
This operation does not change the homotopy type of $\Sigma$ nor the boundary map $\partial f:\partial\Sigma\rightarrow\coprod_{i}S^{1}$. It also preserves the class $f_{*}[\Sigma]=n(\Sigma)\alpha$, as well as monotonicity of $f$. Moreover, it makes the number of cellular discs of $\Sigma$ decrease strictly. Hence, after repeating a finite number of times, we may assume that each connected component of $\Sigma$ contains cellular discs of only one type: either orientation-preserving or orientation-reversing.
@@
Let $S$ be a cellulated oriented compact connected surface. A transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ is said to be non-folded if each connected component of $\Sigma$ contains cellular discs that are either all orientation-preserving or all orientation-reversing.
This terminology is inspired by the picture of Figure 3: when a transverse admissible surface has two adjacent cellular discs of opposite orientations, then it folds those cellular discs onto one another.
@@
Let $\alpha\in H_{2}(S,c)$ be a fixed class. Given a transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, there is a non-folded admissible surface $f^{\prime}:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow\left(S,c\right)$, with $f^{\prime}_{*}\left[\Sigma^{\prime}\right]=n\left(\Sigma^{\prime}\right)\alpha$ such that
@@
Moreover, $f^{\prime}$ can be assumed to be monotone if $f$ was.∎
4.d. Asymptotic promotion to orientation-perfect surfaces
For our purpose, non-folded admissible surfaces will not be enough. We would like admissible surfaces that satisfy the following:
@@
Let $S$ be a cellulated, oriented, compact, connected surface. A transverse admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ is said to be orientation-perfect if there are no two cellular discs in $\Sigma$ that map to the same $2$-cell of $S$ with opposite orientations.
@@
We start with a non-folded, monotone, disc-free, transverse, incompressible, admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, and we assume that $f$ is not orientation-perfect. Fix a small $\varepsilon>0$, and pick a large $N\in\mathbb{N}_{\geq 1}$ such that $\frac{1}{N}\leq\varepsilon$. Let ${\Sigma_{0}}\rightarrow\Sigma$ be a degree-$N$ covering under which the preimage of every connected component of $\Sigma$ is connected. The composite map ${\Sigma_{0}}\rightarrow\Sigma\rightarrow S$ is also a non-folded monotone, transverse, incompressible admissible surface for $\operatorname{scl}(c)$ or $\left\|\alpha\right\|_{\textnormal{Grom}}$, with $\chi^{-}\left({\Sigma_{0}}\right)=N\chi^{-}(\Sigma)$ and $n\left({\Sigma_{0}}\right)=Nn\left(\Sigma\right)$. Since $f$ is non-folded but not orientation-perfect, there are two cellular discs in distinct components of ${\Sigma_{0}}$ that map to the same $2$-cell of $S$ with opposite orientations. We remove those two discs and glue the resulting boundary components to one another in a way that is compatible with the map $f$. The map $f^{\prime}:\left({\Sigma}_{0}^{\prime},\partial{\Sigma}_{0}^{\prime}\right)\rightarrow(S,c)$ resulting from this operation is still a transverse monotone admissible surface; it satisfies $f^{\prime}_{*}\left[{\Sigma}_{0}^{\prime}\right]=Nn\left(\Sigma\right)\alpha$, and
@@
We can then perform the process of §4.c again to ensure that ${\Sigma}_{0}^{\prime}$ is non-folded.
@@
Let $\alpha\in H_{2}(S,c)$ be a fixed class. Given an $\varepsilon>0$ and a transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, there is an orientation-perfect admissible surface $f^{\prime}:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow\left(S,c\right)$ with $f^{\prime}_{*}\left[\Sigma^{\prime}\right]=n\left(\Sigma^{\prime}\right)\alpha$ such that
@@
Moreover, $f^{\prime}$ can be assumed to be monotone if $f$ was.∎
@@
4.e. Connectedness of links
We are going to perform one more modification on our admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$: we want to ensure that links of vertex discs in $\Sigma$ are connected.
Consider the $2$-complex $\bar{\Sigma}$ obtained from $\Sigma$ by collapsing all vertex discs to vertices and all $1$-handles to edges — hence, $f$ induces a combinatorial map $\bar{f}:\bar{\Sigma}\rightarrow S$, but $\bar{\Sigma}$ may not be a surface.
Definition 4.7.
Assume first that $v$ lies in the interior of $S$; an example is depicted in Figure 5.
[…]
Figure 5. A vertex disc with disconnected link mapping to a vertex in the interior of $S$.
The main point is that there is a $2$-cell in $S$ between any two consecutive edges around $v$; in particular, there is a $2$-cell $\sigma$ that lies between two successive edges whose preimages are $1$-handles with one end each on $\partial\Sigma$, as in Figure 5. (Note that one may have to pick a sequence of $2$-cells, rather than a single $2$-cell, but the argument works in the same way.) Now we perform a homotopy that moves the image of $\partial\Sigma$ across the $2$-cell $\sigma$: see Figure 6.
[…]
Figure 6. Making links of vertex discs connected (interior case).
This operation decreases the number of connected components in the link of $D$ (or more precisely, of its image in $\bar{\Sigma}$). Note that new vertex discs may have been created (such as the one mapping to $u_{5}$ in Figure 6), but they all have connected link. As for preexisting vertex discs of $\Sigma$, the operation doesn’t impact the number of connected components of their links. Hence, if $\left\{D_{i}\right\}_{i}$ is the set of vertex discs of $\Sigma$ mapping to the interior of $S$, and $k_{i}$ denotes the number of connected components of the link of $D_{i}$, then we have made the quantity $\sum_{i}\left(k_{i}-1\right)$ decrease strictly. We may therefore iterate to ensure that all vertex discs mapping to the interior of $S$ have connected link.
We deal with the case where $v$ lies on the boundary in the following way. We thicken $S$ by gluing a cellulated annulus to each of its boundary components. This modifies the cellular structure of $S$ but not its homeomorphism type (in other words, $S$ has been replaced by another cell complex $S^{\prime}$ with more cells, but with $S^{\prime}$ homeomorphic to $S$), and this preserves all the properties of the map $f$ — in particular, $f$ is transverse for the new cellulation of $S$, and its orientation properties are unchanged as the images of preexisting $2$-cells are not modified. Now all of the vertex discs with disconnected link map to the interior of $S$. Therefore, we can apply the operation described above to make all their links connected. This may create some new vertex discs mapping to the boundary, but they will have connected link. Hence, after both operations, all vertex discs have connected link.
As this process creates new cellular discs in $\Sigma$ (because of the treatment of vertex discs mapping to the interior), we need to check that the non-folding and the orientation-perfect properties are preserved. This is easily seen to be true for the non-folding property: whenever we create a new cellular disc, it maps to $S$ with the same orientation as neighbouring cellular discs, and hence as its entire connected component in $\Sigma$. However, this may not be the case for the orientation-perfect property: it may be that a new cellular disc maps to a $2$-cell $\sigma$ with a different orientation from other cellular discs in the preimage of $\sigma$. If this happens, then we apply the asymptotic promotion process of §4.d again. This decreases the number of connected components of $\Sigma$ strictly. In doing this, we may have created new vertices with disconnected link, but we can apply the entire process again to make their link connected. To sum up, what we are doing here is alternately making all links connected, then applying the asymptotic promotion operation to ensure that the admissible surface is still orientation-perfect. A key observation is that the total number of applications of the asymptotic promotion process is bounded by the initial number of connected components of $\Sigma$ (as we never create any new component). This guarantees on the one hand that the process terminates, and on the other hand that we can control the error $\varepsilon$ that is created by asymptotic promotion (see Lemma 4.5).
We therefore obtain the following:
Lemma 4.8 (Admissible surfaces with connected links).
Let $\alpha\in H_{2}(S,c)$, and let $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ be a transverse incompressible admissible surface with $f_{*}[\Sigma]=n(\Sigma)\alpha$.
  1. (i)
    There is a $2$-complex $S^{\prime}$ homeomorphic to $S$, and a non-folded admissible surface $f:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow(S^{\prime},c)$ that has connected links and such that $-\chi^{-}\left(\Sigma^{\prime}\right)/n\left(\Sigma^{\prime}\right)\leq-\chi^{-}(\Sigma)/n(\Sigma)$.
  2. (ii)
    For any $\varepsilon>0$, there is a $2$-complex $S^{\prime}$ homeomorphic to $S$, and an orientation-perfect admissible surface $f:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow(S^{\prime},c)$ that has connected links and such that $-\chi^{-}\left(\Sigma^{\prime}\right)/n\left(\Sigma^{\prime}\right)\leq-\chi^{-}(\Sigma)/n(\Sigma)+\varepsilon$.
Moreover, $f^{\prime}$ can be assumed to be monotone if $f$ was.∎
@@
Let $S$ be an oriented, compact, connected surface, let $c\in C_{1}\left(\pi_{1}S;\mathbb{Z}\right)$ be an integral chain, and $\alpha\in H_{2}(S,c;\mathbb{Q})$. Then
@@
    where the infimum is taken over all admissible surfaces $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ that are orientation-perfect, monotone, disc-free, transverse, incompressible, with connected links for some cellulation of $S$.
@@
    If there exists an extremal surface for $\left\|\alpha\right\|_{\textnormal{Grom}}$ (i.e. realising the infimum in Definition 2.10), then there exists one which is non-folded, monotone, disc-free, transverse, incompressible, with connected links for some cellulation of $S$.
@@
This follows from §4.b (for transversality) and Lemmas 4.1 (for monotonicity), 4.3 (for the non-folding property), 4.5 (for the orientation-perfect property), and 4.8 (for connected links). ∎

### v3

@@ Acknowledgements
First and foremost, I would like to thank my supervisor Henry Wilton for our weekly discussions and his constant support. Thanks are also due to Lvzhou Chen, Francesco Fournier-Facio, and Kevin Li for helpful conversations, as well as to the anonymous referee for careful reading and detailed comments. Financial support from an EPSRC PhD studentship is gratefully acknowledged.
@@ 2. Scl and the relative Gromov seminorm
In Definition 2.3, replacing $\mathbb{R}$ with $\mathbb{Q}$ or $\mathbb{Z}$ does not change what it means for a group homomorphism to be $\operatorname{scl}$-preserving or a (strong) isometric embedding for $\operatorname{scl}$.
@@
    with $k_{i}\in\mathbb{Z}\smallsetminus\{0\}$. Then $\varphi$ is a strong isometric embedding for $\operatorname{scl}$.
@@ 4. Standard form for admissible surfaces
Observe first that we can harmlessly remove any disc- or sphere-component of $\Sigma$, since this does not change $\chi^{-}(\Sigma)$. We then say that $\Sigma$ is disc- and sphere-free.
@@
In the context of $\operatorname{scl}$, it is a classical fact [cal-scl]*Proposition 2.13 that one can work with monotone admissible surfaces only. Our definition of monotonicity allows us to adapt this to the relative Gromov seminorm [^4] This would be false with the usual definition — see Footnote 3.. Calegari’s proof [cal-scl]*Proposition 2.13 works in our context, where one should deal with boundary components of $\Sigma$ mapping to distinct components of $\coprod_{i}S^{1}$ separately. See [m:phd]*Lemma II.1.4 for more details.
@@
Consider an admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$ in $H_{2}(S,c)$. We may assume that $f$ is transverse, incompressible, monotone, and disc- and sphere-free, as explained in §4.a and §4.b.
We now want to use the fact that $S$ is a surface to ensure that $\Sigma$ is ‘thick enough’, in the sense that its vertex discs have connected links.
More precisely, consider the $2$-complex $\bar{\Sigma}$ obtained from $\Sigma$ by collapsing all vertex discs to vertices and all $1$-handles to edges — hence, $f$ induces a combinatorial map $\bar{f}:\bar{\Sigma}\rightarrow S$, but $\bar{\Sigma}$ may not be a surface (see for instance Figure 3: the vertex disc at the centre of $\Sigma$ becomes a disconnecting vertex in $\bar{\Sigma}$).
Definition 4.2.
Assume first that $v$ lies in the interior of $S$; an example is depicted in Figure 3 (all the $2$-cells on the picture are triangles for simplicity, but the general case is similar).
[…]
Figure 3. A vertex disc with disconnected link mapping to a vertex in the interior of $S$ (note that the map $f$ is orientation-preserving on the blue cellular disc but orientation-reversing on the green one).
The main point is that there is a $2$-cell in $S$ between any two consecutive edges around $v$; in particular, there is a sequence of $2$-cells $\sigma_{1},\dots,\sigma_{\ell}$ lying between two successive edges whose preimages are $1$-handles with one end each on $\partial\Sigma$, as in Figure 3. Now we perform a homotopy that moves the image of $\partial\Sigma$ across the $2$-cells $\sigma_{1},\dots,\sigma_{\ell}$: see Figure 4.
[…]
Figure 4. Making links of vertex discs connected (interior case).
The new map $f:\Sigma\rightarrow S$ defines an admissible surface $\left(\Sigma,\partial\Sigma\right)\rightarrow\left(S,c\right)$. Note that $f$ has been modified by a homotopy, so we still have $f_{*}[\Sigma]=n(\Sigma)\alpha$.
This operation decreases the number of connected components in the link of $D$ (or more precisely, of its image in $\bar{\Sigma}$). Note that new vertex discs may have been created (such as the one mapping to $u_{4}$ in Figure 4), but they all have connected link. As for preexisting vertex discs of $\Sigma$, the operation doesn’t impact the number of connected components of their links. Hence, if $\left\{D_{i}\right\}_{i}$ is the set of vertex discs of $\Sigma$ mapping to the interior of $S$, and $k_{i}$ denotes the number of connected components of the link of $D_{i}$, then we have made the quantity $\sum_{i}\left(k_{i}-1\right)$ decrease strictly. We may therefore iterate to ensure that all vertex discs mapping to the interior of $S$ have connected link.
We deal with the case where $v$ lies on the boundary in the following way. We thicken $S$ by gluing a cellulated annulus to each of its boundary components. This modifies the cellular structure of $S$ but not its homeomorphism type (in other words, $S$ has been replaced with another cell complex $S^{\prime}$ with more cells, but with $S^{\prime}$ homeomorphic to $S$), and this preserves all the properties of the map $f$ — in particular, $f$ is transverse for the new cellulation of $S$. Now all of the vertex discs with disconnected link map to the interior of $S$. Therefore, we can apply the operation described above to make all their links connected. This may create some new vertex discs mapping to the boundary, but they will have connected link. Hence, after both operations, all vertex discs have connected link.
We therefore obtain the following:
Lemma 4.3 (Admissible surfaces with connected links).
Fix $\alpha\in H_{2}\left(S,c\right)$, and let $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ be a transverse incompressible admissible surface with $f_{*}[\Sigma]=n(\Sigma)\alpha$. Then, after possibly changing the cellular structure on $S$, the map $f$ may be homotoped to a transverse incompressible admissible surface with connected links.∎
Note that since our admissible surface was modified by a homotopy, the properties of being incompressible, monotone, and disc- and sphere-free are preserved.
4.d. Folding
@@
The key properties of admissible surfaces that we will need in the surface group case are related to orientation. Indeed, both $S$ and the admissible surface $\Sigma$ are oriented. Since $f$ is transverse and cellular discs map homeomorphically into $S$, they can be of two types: either they preserve the orientation or they reverse it. Having cellular discs of opposite orientations is undesirable, and we are now going to modify $\Sigma$ to avoid this situation as much as possible.
We assume that $f:\left(\Sigma,\partial\Sigma\right)\rightarrow\left(S,c\right)$ is transverse, incompressible, monotone, disc- and sphere-free, and with connected links.
Suppose first that there is a connected component of $\Sigma$ containing cellular discs of two different types — i.e. one is orientation-preserving with respect to $f$ and the other is orientation-reversing. Since $f:\left(\Sigma,\partial\Sigma\right)\rightarrow\left(S,c\right)$ has connected links, any two cellular discs in $\Sigma$ mapping to cells in $S$ with a common vertex on their boundaries must be connected by a path of cellular discs and $1$-handles in $\Sigma$. It follows that any two cellular discs in the same connected component of $\Sigma$ are connected by a path of cellular discs and $1$-handles.
@@
Therefore, $\Sigma$ must contain two cellular discs of opposite types that are adjacent via a $1$-handle. Since $S$ is a surface, those two cellular discs must map to the same $2$-cell of $S$, and we are in the situation of Figure 5 (pictures are given for the case of a $2$-cell of degree $3$, but the general case is similar) — in other words, $f$ folds those two cellular discs onto one another.
@@
This operation does not change the homotopy type of $\Sigma$ nor the boundary map $\partial f:\partial\Sigma\rightarrow\coprod_{i}S^{1}$. It also preserves the class $f_{*}[\Sigma]=n(\Sigma)\alpha$, as well as transversality, incompressibility, monotonicity of $f$, and the property of being disc- and sphere-free. Moreover, it makes the number of cellular discs of $\Sigma$ decrease strictly. Hence, after repeating a finite number of times, we may assume that each connected component of $\Sigma$ contains cellular discs of only one type: either orientation-preserving or orientation-reversing.
@@
Let $S$ be a cellulated, oriented, compact, connected surface. A transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ is said to be non-folded if each connected component of $\Sigma$ contains cellular discs that are either all orientation-preserving or all orientation-reversing.
However, the operation just described could have created some vertex discs with disconnected link in $\Sigma$ (as removing cellular discs amounts to deleting edges in the links of vertex discs).
To fix this, we would like to successively apply the operation described above and the one of §4.c. This relies on the following crucial observation: in the process described in §4.c, we can always choose the orientation of the cellular discs that we add. Indeed, the added cellular discs correspond to a path in the link of $v$ in $S$, and this link is a circle, so there are two possible paths, one corresponding to adding cellular discs of positive orientation to $\Sigma$, and the other corresponding to adding cellular discs of negative orientation. For example, Figure 7 shows two possible choices that make the link of the vertex disc of Figure 3 connected. In particular, when applying the operation of §4.c, we can assume that we are only adding cellular discs of positive orientation.
[…]
Figure 7. Two possible choices for making the link of the vertex disc of Figure 3 connected.
We can now apply the following procedure: we alternately apply the operation of §4.c to make links connected — and with the choice of only adding cellular discs of positive orientation — and then the operation described above to remove folding. Each iteration of the latter removes one cellular disc of each orientation, and each iteration of the former does not increase the number of cellular discs of negative orientation. Hence, the total number of discs of negative orientation decreases strictly at each pair of iterations, ensuring that the process terminates in an admissible surface that is both non-folded and with connected links.
Note that the operations just described do not impact the boundary map (up to homotopy), and hence the monotonicity, of $f$. Hence we can first apply Lemma 4.1 and ensure that the resulting surface is monotone. We can also readily remove any disc- or sphere-component.
This proves the following:
@@
Fix $\alpha\in H_{2}\left(S,c\right)$. Given a transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, there is a non-folded admissible surface $f^{\prime}:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow\left(S,c\right)$ with connected links, with $f^{\prime}_{*}\left[\Sigma^{\prime}\right]=n\left(\Sigma^{\prime}\right)\alpha$, and such that
@@
Moreover, $f^{\prime}$ can be assumed to be monotone and disc- and sphere-free.∎
4.e. Asymptotic promotion to orientation-perfect surfaces
In order to obtain isometric embedding results for surface groups in §5, we will need admissible surfaces to satisfy the following orientation property:
@@
Let $S$ be a cellulated, oriented, compact, connected surface. A transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,\gamma)$ is orientation-perfect if there are no two cellular discs in $\Sigma$ that map to the same $2$-cell of $S$ with opposite orientations.
@@
We start with a transverse, incompressible, monotone, disc- and sphere-free, non-folded admissible surface with connected links $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, and we assume that $f$ is not orientation-perfect. Fix a small $\varepsilon>0$, and pick a large $N\in\mathbb{N}_{\geq 1}$ such that $\frac{1}{N}\leq\varepsilon$. Let ${\Sigma_{0}}\rightarrow\Sigma$ be a degree-$N$ covering under which the preimage of every connected component of $\Sigma$ is connected. The composite map ${\Sigma_{0}}\rightarrow\Sigma\rightarrow S$ is also a transverse, incompressible, monotone, disc- and sphere-free, non-folded admissible surface with connected links, with $\chi^{-}\left({\Sigma_{0}}\right)=N\chi^{-}(\Sigma)$ and $n\left({\Sigma_{0}}\right)=Nn\left(\Sigma\right)$. Since $f$ is non-folded but not orientation-perfect, there are two cellular discs in distinct components of ${\Sigma_{0}}$ that map to the same $2$-cell of $S$ with opposite orientations. We remove those two discs and glue the resulting boundary components to one another in a way that is compatible with the map $f$. There is an admissible surface $f^{\prime}:\left({\Sigma}_{0}^{\prime},\partial{\Sigma}_{0}^{\prime}\right)\rightarrow(S,c)$ resulting from this operation, which is still transverse, incompressible, monotone, disc- and sphere-free; it satisfies $f^{\prime}_{*}\left[{\Sigma}_{0}^{\prime}\right]=Nn\left(\Sigma\right)\alpha$, and
@@
We can then perform the process of §4.d again to ensure that ${\Sigma}_{0}^{\prime}$ is non-folded and with connected links.
@@
Fix $\alpha\in H_{2}\left(S,c\right)$. Given an $\varepsilon>0$ and a transverse incompressible admissible surface $f:(\Sigma,\partial\Sigma)\rightarrow(S,c)$ with $f_{*}[\Sigma]=n(\Sigma)\alpha$, there is an orientation-perfect admissible surface $f^{\prime}:\left(\Sigma^{\prime},\partial\Sigma^{\prime}\right)\rightarrow\left(S,c\right)$, with $f^{\prime}_{*}\left[\Sigma^{\prime}\right]=n\left(\Sigma^{\prime}\right)\alpha$, and such that
@@
Moreover, $f^{\prime}$ can be assumed to be monotone, disc- and sphere-free, non-folded, and with connected links.∎
@@
Let $S$ be an oriented, compact, connected surface, let $c\in C_{1}\left(\pi_{1}S;\mathbb{Z}\right)$ be an integral chain, and $\alpha\in H_{2}\left(S,c;\mathbb{Q}\right)$. Then
@@
    where the infimum is taken over all admissible surfaces $f:(\Sigma,\partial\Sigma)\rightarrow\left(S,c\right)$ that are transverse, incompressible, monotone, disc- and sphere-free, orientation-perfect, with connected links for some cellulation of $S$.
@@
    If there exists an extremal surface for $\left\|\alpha\right\|_{\textnormal{Grom}}$ (i.e. realising the infimum in Definition 2.10), then there exists one which is transverse, incompressible, monotone, disc- and sphere-free, non-folded, with connected links for some cellulation of $S$.
@@
This follows from §4.b (for transversality) and Lemmas 4.1 (for monotonicity), 4.3 (for connected links), 4.5 (for the non-folding property), and 4.7 (for the orientation-perfect property). ∎
Remark 4.10.
By the discussion of §4.d, an orientation-perfect admissible surface is automatically non-folded. It follows that an admissible surface in perfect standard form is also in standard form.